# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

### v2

@@ 1. Introduction
  • Based on the characterization, we prove the sparse Moment-SOS hierarchy is tight when some classical sufficiency optimality conditions hold or when each individual equality constraining variety is finite. In particular, when the RIP holds, we show that this sparse hierarchy is tight if each individual constraining variety is finite.
@@ 2.1. Ideals and SOS polynomials
Let $K$ be the set as in (2.5). Then the sparse Positivstellensatz (2.6) holds if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
@@ 4.1. The convex case
Suppose that each $h_{i}$ is linear (or it does not appear), all $f_{i}$ and $-g_{ij}$ are convex, and $u$ is a minimizer of (1.1). If (4.3) holds, then Assumption 4.1 holds. Furthermore, if in addition, the sub-Hessian
@@
$p_{i}(x)\coloneqq-(x-u)^{T}\big{[}\nabla f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i}(u_{\Delta_{i}})-\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i}(u_{\Delta_{i}})\big{]}-f_{i}(u_{\Delta_{i}}).$
@@
$\nabla f_{i}(u_{\Delta_{i}})+\nabla p_{i}(u_{\Delta_{i}})=\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i}(u_{\Delta_{i}})+\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i}(u_{\Delta_{i}}),$
@@
Assume the feasible set $K\neq\emptyset$ and the minimum value $f_{\min}>-\infty$. Suppose each $h_{i}$ is linear or it does not appear, each $f_{i}$ and $-g_{ij}$ are SOS-convex in $x_{\Delta_{i}}$. Then, we have:
@@ 4.2. The case of sufficient optimality conditions
Under the archimedeanness, the dense Moment-SOS hierarchy in [16] is tight when linear independence constraint qualification condition (LICQC), strict complementarity condition (SCC) and second order sufficient condition (SOSC) hold at every minimizer, as shown in [31]. We refer to [32, Chapter 5.1] for these conditions. For sparse Moment-SOS relaxations, we have a similar conclusion under Assumption 4.1.
@@ 4.3. The case of finite sets
For a set $S\subset\mathbb{R}^{\Gamma}$ with $\Gamma\supseteq I$, we denote the projection
@@ 6. Some examples
Consider the convex optimization problem:
@@
Each $f_{i}$ is SOS-convex. For instance, $f_{1}$ is SOS-convex since
$\nabla^{2}f_{1}(x_{1},x_{2})=\begin{bmatrix}30x_{1}^{4}&9x_{1}^{2}x_{2}^{2}\\ 9x_{1}^{2}x_{2}^{2}&30x_{2}^{4}\end{bmatrix}=\begin{bmatrix}x_{1}^{2}&0\\ 0&x_{2}^{2}\end{bmatrix}\begin{bmatrix}30&9\\ 9&30\end{bmatrix}\begin{bmatrix}x_{1}^{2}&0\\ 0&x_{2}^{2}\end{bmatrix}.$
By Theorem 4.3, the sparse relaxation (3.2)-(3.3) is tight for all $k\geq 3$. Solving it with $k=3$, we get $f^{spa}_{3}\approx-2.2561.$ By Theorem 4.3 (ii), we get $f_{\min}=f^{spa}_{3}$ and the minimizer $(-0.6036,0.6852,-0.7092).$
@@ 7. Some proofs
$\hat{y}\coloneqq(\hat{y}_{\alpha})_{\alpha\in\mathbb{U}_{k}}\quad\text{where each}\quad\hat{y}_{\alpha}=\left(x_{\Delta_{i}}^{*}\right)^{\alpha}.$
Let $\tilde{y}\coloneqq(\tilde{y}_{\alpha})_{\alpha\in\mathbb{U}_{k}}$ be the tms such that $y^{*}=\rho\hat{y}+(1-\rho)\tilde{y}$, then
@@
Since $f_{t}^{smo}=f_{k}^{smo}$, the truncation $y^{*}|_{2t}$ is a minimizer for (3.3) with $k=t$, so
@@
$\in$ $\operatorname{Ideal}[h]_{spa,k}+\operatorname{QM}[g]_{spa,k}.$

### v3

@@ 1. Introduction
  • Based on the characterization, we prove the sparse Moment-SOS hierarchy is tight when some classical sufficient optimality conditions hold or when each individual equality constraining variety is finite. In particular, when the RIP holds, we show that this sparse hierarchy is tight if each individual constraining variety is finite.
@@ 2.1. Ideals and SOS polynomials
Let $K$ be the set as in (2.5). Then a sparse version of Positivstellensatz (2.6) holds if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
@@ 4.1. The convex case
Suppose that each $h_{i}$ is linear (or it does not appear), all $f_{i}$ and $-g_{i,j}$ are convex, and $u$ is a minimizer of (1.1). If (4.3) holds, then Assumption 4.1 holds. Furthermore, if in addition, the sub-Hessian
@@
$p_{i}(x)\coloneqq-(x-u)^{T}\big{[}\nabla f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i,j}(u_{\Delta_{i}})-\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i,j}(u_{\Delta_{i}})\big{]}-f_{i}(u_{\Delta_{i}}).$
@@
$\nabla f_{i}(u_{\Delta_{i}})+\nabla p_{i}(u_{\Delta_{i}})=\sum_{j=1}^{s_{i}}\lambda_{i,j}\nabla g_{i,j}(u_{\Delta_{i}})+\sum_{j=1}^{\ell_{i}}\nu_{i,j}\nabla h_{i,j}(u_{\Delta_{i}}),$
@@
Assume the feasible set $K\neq\emptyset$ and the minimum value $f_{\min}>-\infty$. Suppose each $h_{i}$ is linear or it does not appear, each $f_{i}$ and $-g_{i,j}$ are SOS-convex in $x_{\Delta_{i}}$. Then, we have:
@@ 4.2. The case of sufficient optimality conditions
Under the archimedeanness, the dense Moment-SOS hierarchy in [16] is tight when linear independence constraint qualification condition (LICQC), strict complementarity condition (SCC) and second order sufficient condition (SOSC) hold at every minimizer, as shown in [31]. We refer to [32, Section 5.1] for these conditions. For sparse Moment-SOS relaxations, we have a similar conclusion under Assumption 4.1.
@@ 4.3. The case of finite sets
For a set $S\subseteq\mathbb{R}^{\Gamma}$ with $\Gamma\supseteq I$, we denote the projection
@@ 6. Some examples
Consider the optimization problem:
@@
The sparse relaxation (3.2)-(3.3) is tight for all $k\geq 3$. Solving it with $k=3$, we get $f^{spa}_{3}\approx-2.2561.$ We get $f_{\min}=f^{spa}_{3}$ and the minimizer $(-0.6036,0.6852,-0.7092).$
@@ 7. Some proofs
$\hat{y}\coloneqq(\hat{y}_{\alpha})_{\alpha\in\mathbb{U}_{t}}\quad\text{where each}\quad\hat{y}_{\alpha}=\left(x_{\Delta_{i}}^{*}\right)^{\alpha}.$
Let $\tilde{y}\coloneqq(\tilde{y}_{\alpha})_{\alpha\in\mathbb{U}_{t}}$ be the tms such that $y^{*}=\rho\hat{y}+(1-\rho)\tilde{y}$, then
@@
Since $f_{t}^{smo}=f_{k}^{smo}$, the truncation $y^{*}|_{spa,2t}$ is a minimizer for (3.3) with $k=t$, so
@@
$\in$ $\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.$